OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^2+2x+y^2+4y-11=0

OpenStudy (aravindg):

the general equation of a circle is \[x^2+y^2+2gx+2fy+c=0\]

OpenStudy (aravindg):

where (-g,-f) is the centre

OpenStudy (aravindg):

*center

OpenStudy (dumbcow):

solution is posted on your other post

OpenStudy (anonymous):

I got the center as 1,2 and the radius 16

OpenStudy (aravindg):

radius=\(\sqrt{g^2+f^2-c}\)

OpenStudy (aravindg):

Seems @dumbcow already answered your question :)

OpenStudy (anonymous):

maybe easier to write in the form \[(x-h)^2+(y-k)^2=r^2\] get the center is \((h,k)\) and radius is \(r\) do you know how to complete the square?

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[x^2+2x+y^2+4y-11=0\] \[x^2+2x+y^2+4y=11\] \[(x+1)^2+(y+2)^2=11+1^2+2^2=16\]

OpenStudy (anonymous):

i got the center as (1,2) and radius 16

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

the form is \[(x-h)^2+(y-k)^2=r^2\] and you have \[(x+1)^2+(y+2)^2=16\]

OpenStudy (anonymous):

oh! (2,1)

OpenStudy (anonymous):

that means the center is \((-1,-2)\) and the radius is \(4\)

OpenStudy (anonymous):

no, not (2,1) but rather (-1,-2)

OpenStudy (anonymous):

Oh i see my mistake thank you

OpenStudy (dumbcow):

here is a good reference to practice and learn completing the square http://www.purplemath.com/modules/sqrcircle.htm